OpenStudy (anonymous):

Write the equation of the circle containing the points J(-3, -3), K(1, -3), and L(1, 1). Show all work to receive credit.

OpenStudy (anonymous):

Set up a system of equations. (-3 - h)^2 + (-3 - k)^2 = r^2 (1 - h)^2 + (-3 - k)^2 = r^2 (1 - h)^2 + (1 - k)^2 = r^2 Start with the first two: (-3 - h)^2 + (-3 - k)^2 = r^2 -1((1 - h)^2 + (-3 - k)^2 = r^2) (-3 - h)^2 + (-3 - k)^2 = r^2 -(1 - h)^2 - (-3 - k)^2 = -r^2 (-3 - h)^2 - (1 - h)^2 = 0 9 + 6h + h^2 - (1 - 2h + h^2) = 0 9 + 6h + h^2 - 1 + 2h - h^2 = 0 8h + 8 = 0 h = -1 Now go into the second pair of equations: (1 - h)^2 + (-3 - k)^2 = r^2 -1((1 - h)^2 + (1 - k)^2 = r^2) (1 - h)^2 + (-3 - k)^2 = r^2 -(1 - h)^2 - (1 - k)^2 = -r^2 (-3 - k)^2 - (1 - k)^2 = 0 9 + 6k + k^2 - (1 - 2k + k^2) = 0 9 + 6k + k^2 - 1 + 2k - k^2 = 0 8k + 8 = 0 k = -1 Use the distance formula. Center: (-1, -1) r = 2√2 Rewrite the formula (x - h)^2 + (y - k)^2 = r^2 (x - (-1))^2 + (y - (-1))^2 = 2√2^2 (x + 1)^2 + (y + 1)^2 = 8

OpenStudy (anonymous):

Man. That took forever! It should take less time when not showing every little step. :)

OpenStudy (anonymous):

ok. wow. thank you so much. I had no idea what i was doing. this helps alot!

OpenStudy (anonymous):

You're welcome :)

OpenStudy (amistre64):

:) i just did that for another problem; glad to see i was in the right track tho